OpenStudy (anonymous):

Suppose the polynomial function below represents the power generated by a wind turbine, where x represents the wind speed in meters per second and y represents the kilowatts generated. Interpret ƒ(10). ƒ(x) = 0.08x3 + x2 + x + 0.26

OpenStudy (campbell_st):

so substitute x = 10 so you will be finding the kilowatts generated when the wind speed is 10 metres per sec.

OpenStudy (anonymous):

ƒ(x) = 0.08x^3 + x^2 + x + 0.26 = (0.08 *10^3) + 10^2 + 10 + 0.26 =80+10^2+10+.26 =80+100+10+.26 =190.26 Kilowatts